OpenStudy (anonymous):

Help : : : Question from numerical analysis. Give an example with proper justification to show that the condition of convergence of fixed point iteration for numerical solution of equations is sufficient and by no means necessary.

OpenStudy (anonymous):

@amistre64 @ganeshie8 @eliassaab

OpenStudy (amistre64):

i must say that this is not going to be workable to me. i simple do not have the proper thrms and or justifications (vocabulary) to assess it.

OpenStudy (anonymous):

You just give me the example @amistre64 and a brief outline, i can manage the rest.:)

OpenStudy (anonymous):

@gorv

OpenStudy (gorv):

it will be too long

OpenStudy (gorv):

bt i can give u idea

OpenStudy (gorv):

like u have to iterate ant function

OpenStudy (anonymous):

Yeah thats what i need.:)

OpenStudy (gorv):

when u itrate let the ans is 0.2345098655

OpenStudy (gorv):

then keep iterating

OpenStudy (gorv):

at one point when the ans will come like ....four point after decimal are similar thn that will be your ans

OpenStudy (gorv):

@Princer_Jones u got me ??

OpenStudy (anonymous):

@gorv

OpenStudy (gorv):

ohh u have any function??

OpenStudy (gorv):

we first find \[\phi(x)\]

OpenStudy (anonymous):

no thats what we need to construct , by the question. If i already had the function, then i would have completed solving it.:)

OpenStudy (gorv):

such that it derivative is less than one

OpenStudy (anonymous):

Derivative will be grater than 1... if it is less than 1, then it will always converge, since the condition is sufficient

OpenStudy (gorv):

let x^4-x-10=0

OpenStudy (gorv):

ok should we proceed??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

in fixed point iteration , the equation becomes x=x^4-10 Hence here fi(x)=x^4-10 whose derivative is 4x^3

OpenStudy (gorv):

whic is greater than 1

OpenStudy (gorv):

x^4-x=10

OpenStudy (gorv):

x(x^3-1)=10

OpenStudy (gorv):

x=10/(x^3-1)

OpenStudy (gorv):

now??

OpenStudy (anonymous):

oh you took different expression for fi(x)?.

OpenStudy (anonymous):

Here fi(x)=10/(x^3-1)?

OpenStudy (gorv):

yeah..and its derivative will be <1

OpenStudy (gorv):

that how we select the function ..ok gotcha??

OpenStudy (anonymous):

hmm but we need the derivative to be greater than 1 and show that it still converges...

OpenStudy (gorv):

u can go with any condition

OpenStudy (gorv):

this one will be smallest

OpenStudy (gorv):

u can go with x=x^4-10

OpenStudy (gorv):

but it will take longer

OpenStudy (anonymous):

The condition of convergence is that the derivative is less than 1. But we have to construct a function such that the derivative is greater than 1 , but it still converges, and then only it will prove that going against the condition of convergence too we have a solution, that means the condition is sufficient and no longer necessary

OpenStudy (gorv):

well thn u can go with your function

OpenStudy (anonymous):

x^4-10?

OpenStudy (anonymous):

@ganeshie8 help please//